OpenStudy (zackattack350):

Please help i dont understand. if y varies inversely as x find the constant of variation k and the inverse variation equation for the situation y=7 when x=2

OpenStudy (anonymous):

y=k/x

OpenStudy (anonymous):

This is the equation showing y is inverse to x.

OpenStudy (anonymous):

Plug in the numbers gets you...

OpenStudy (zackattack350):

do i just plug it in? and that will be the answer for the inverse equation? and the constant of K will be the answer?

OpenStudy (anonymous):

k=14!

OpenStudy (anonymous):

yeah just plug it in. I cant do math lol

OpenStudy (zackattack350):

k=7/2

OpenStudy (zackattack350):

so... 3.5

OpenStudy (anonymous):

no y=k/x k=y*x

OpenStudy (anonymous):

7*2 is 14

OpenStudy (zackattack350):

Oh. so y=k/x is inverse

OpenStudy (zackattack350):

so its 14

OpenStudy (anonymous):

yesh :3

OpenStudy (zackattack350):

thanks!